OpenStudy (anonymous):

find two choices for t such that the distance between (3,-2) and (1,t) equals 5

OpenStudy (marigirl):

What I would do is use the distance formula \[s=\sqrt{(x _{2}}-x _{1})^2+(y _{2}-y _{1})^2\] we are given two coordinates (3, -2) and (1,t) \[(x _{1},x _{2})(y _{1},y _{2})\] lets sub it in: \[5=\sqrt{(-2-3)+(t-1)}\] \[5=\sqrt{-5+(t-1)}\] \[5=\sqrt{-6+t}\] Now solve for t